Graph the function. X+1 17) f(x)=1-8 [-X+4 if -9<x<3 if x=3 fx>3 HHHHHHHHH> - 10 -5 5 10 X HiiHiHitta
Answer #1

-g<x<3 x= 3 ze > D3 ut f = 9 -8 if 1-x+4 if 66-9): -9H=18 H1-5) = -541-24 $(2) = 2+1:3 f(-1) = a1415-0 (-9,8 fb.h) = 4th +(3t